Grade 11Physical Chemistry

2NaCl+H2SO4-----> Na2SO4+2HCl....what mass of 80% pure Na2SO4 can be obtained from 200 kg 90% pure salt?

Harsh
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
let x be the mass of Na2SO4
80% of x => 0.8x
moles = 0.8x/142
 
200 * 1000gm
pure salt =  (90/100)*200000 = 180000gm
moles = 180000/58
from eq. concept.
moles * valence factor = constant.
180000/58 * 1 = 0.8x/142 * 2
calculate x from here.
 
U can also do it by mole concept – 1.